# Automorphism group schemes of lattice vertex operator algebras

### v1

@@ 3.3. Action of orthogonal transformations on the lattice vertex algebra
The fact that this is a faithful group action follows from 3.3.2. The fact that the action is by vertex operator algebra automorphisms follows from the multiplication formulas for elements of $V_{L,R}$ in the case $L$ is the Leech lattice and $R$ is a $\mathbb{Q}$-algebra, this is [Frenkel-Lepowsky-Meurman 1988] Theorem 10.4.6, but the proof does not depend on the particular lattice. When $R=\mathbb{Z}$, the formulas follow from base change to $\mathbb{Q}$, and the proof for general $R$ follows by base change from $\mathbb{Z}$. ∎
@@
The action of $O(\tilde{L})$ on $V_{L}$ normalizes the $T_{L}$-action.
@@ 4.2. Local Conjugation
We wish to show that $\operatorname{Aut}V_{L}$ is a product of $G_{L}$ and $O(\tilde{L})$. The following lemma lets us reduce this problem to more local questions.
@@
To show that $\operatorname{Aut}V_{L}$ is a product of $G_{L}$ and $O(\tilde{L})$, it suffices to show that, given a commutative ring $S$ and an $S$-point $g$ of $\operatorname{Aut}V_{L}$, there is a faithfully flat cover $\{U_{i}\to\operatorname{Spec}S\}$ such that each $g|_{U_{i}}$ lies in the image of the product maps on $U_{i}$.
@@
By Proposition 4.1.2, $G_{L}$ is normal in $\operatorname{Aut}V_{L}$, so conjugation by $g$ takes $T_{L,S}$ to some torus $T^{\prime}$ in $G_{L,S}$. By [SGA3] Exp. XXII Corollary 5.3.11, any two maximal tori (of an fpqc locally split fiberwise connected reductive group) are conjugate étale locally on $S$. Thus, there exists a choice of étale cover $\{S\to S_{i}\}$ and $\{g_{i}\in G_{L}(S_{i})\}$ so that conjugation by each $g_{i}$ takes $T^{\prime}_{S_{i}}$ to $T_{L,S_{i}}$. ∎
@@ 4.3. Normalizer of the distinguished torus
Consider the homomorphism $N_{\operatorname{Aut}V_{L}}T_{L}\to\operatorname{Aut}T_{L}$ induced by conjugation. We let $c:N_{\operatorname{Aut}V_{L}}T_{L}\to GL(L)$ be the composite of this homomorphism with the isomorphism $\operatorname{Aut}T_{L}\overset{\sim}{\to}GL_{(}L)$.
@@
The orbit of $T_{L}$ acting on any nonzero vector in $\pi^{L}_{\lambda,R}$ is one dimensional, and contained in a line. Let $u\in\pi^{L}_{\lambda,R}$, and write $\sigma u=\sum_{i\in I}u_{i}$ for $u_{i}\in\pi^{L}_{\mu_{i},R}$, where the map $i\mapsto\mu_{i}$ is injective. If $I$ has more than one element, then the orbit of $T_{L}$ on $\sigma u$ is not contained in a line. This contradicts the fact that $\sigma$ acts linearly and bijectively. We conclude that $I$ is a singleton. ∎
@@
The normalizer of $T_{L}$ in $\operatorname{Aut}V_{L}$ is the product $O(\tilde{L})C_{\operatorname{Aut}V_{L}}T_{L}$. In particular, $c$ induces an isomorphism $(N_{\operatorname{Aut}V_{L}}T_{L})/C_{\operatorname{Aut}V_{L}}T_{L}\to O(L)$ taking a normalizing automorphism to the induced orthogonal transformation on the lattice parametrizing Heisenberg modules in $V_{L}$.
@@
By Proposition 3.3.3, we know that $N_{\operatorname{Aut}V_{L}}T_{L}$ contains $O(\tilde{L})$. Thus, it suffices to show $N_{\operatorname{Aut}V_{L}}T_{L}\subseteq O(\tilde{L})C_{\operatorname{Aut}V_{L}}T_{L}$.
The conjugation homomorphism $c:N_{\operatorname{Aut}V_{L}}T_{L}\to GL(L)$ has kernel $C_{\operatorname{Aut}V_{L}}T_{L}$ and by Lemma 4.3.4, its image lies in $O(L)$. The centralizer of $T_{L}$ in $O(\tilde{L})$ is the kernel of the homomorphism to $O(L)$, and is therefore identified with $\operatorname{Hom}(L,\mu_{2})$. Thus, the image of $O(\tilde{L})$ under $c$ is precisely $O(L)$, and we have an isomorphism. ∎
We can now show that $T_{L}$ is its own centralizer.
@@
It suffices to show that any homogeneous $R$-vertex algebra automorphism $\sigma$ of $V_{L,R}$ fixes $\omega$. By Corollary 4.2.3 and Proposition 4.3.5, we may write $\sigma$ as a product $gh$, where $g\in G_{L}(R^{\prime})$ and $h\in O(\tilde{L})(R^{\prime})$ for $R\hookrightarrow R^{\prime}$tale. The formula for $\omega$ can be written as a sum over dual basis vectors, so it is invariant under orthogonal transformations, and in particular, the action of $O(\tilde{L})$. It therefore suffices to show that $\omega$ is $G_{L}$-invariant, but for this it suffices to show that the action of $T_{L}$ and the infinitesimal actions of root vectors $h_{0}$ are trivial. Because $\omega$ lies in lattice-degree zero, $T_{L}$ acts trivially. For any root vector $h$, the well-known commutator formula $[h_{m},L(n)]=mh_{m+n}$ ([Frenkel-Lepowsky-Meurman 1988] eq. 8.7.20) implies $h_{0}\omega=h_{0}L(-2)\mathbf{1}=L(-2)h_{0}\mathbf{1}=0$. ∎
We now consider the intersection of $O(\tilde{L})$ and $G_{L}$ in $\operatorname{Aut}V_{L}$.
@@
Viewed as a subgroup of $N_{\operatorname{Aut}V_{L}}T_{L}$, $\tilde{W}_{L}$ is a normal subgroup of $O(\tilde{L})$, given by the preimage of $W_{L}$ in $O(L)$ under the $\operatorname{Hom}(L,\mu_{2})$-cover $\pi:O(\tilde{L})\to O(L)$. In particular, the quotient $O(\tilde{L})/\tilde{W}_{L}$ is isomorphic to the outer automorphism group $O(L)/W_{L}$ of $L$.
@@
By Lemma 3.3.4, $O(\tilde{L})$ normalizes $T_{L}$, i.e., we have an embedding $O(\tilde{L})\hookrightarrow N_{\operatorname{Aut}V_{L}}T_{L}$. By Proposition 3.2.2, restriction of $N_{\operatorname{Aut}V_{L}}T_{L}\to N_{\operatorname{Aut}V_{L}}T_{L}/T_{L}=O(L)$ to $O(\tilde{L})$ is a surjective homomorphism to $O(L)$, with kernel $\operatorname{Hom}(L,\mu_{2})$. We therefore have a commutative diagram of extensions, where the vertical arrows are inclusions of normal subgroups, and the left square is a pullback:
[…]
By the universal property of pullback, the pullback along the embeddings $N_{G_{L}}T_{L}\to N_{\operatorname{Aut}V_{L}}T_{L}$ and $O(\tilde{L})\to N_{\operatorname{Aut}V_{L}}T_{L}$ is then given by the diagram:
[…]
where the vertical arrows are inclusions of normal subgroups, and the squares on the top right and bottom left are pullbacks. We obtain an isomorphism $O(\tilde{L})/\tilde{W}_{L}\simeq O(L)/W_{L}$. ∎
@@ 4.4. Main Theorem
Let $L$ be a positive definite even lattice, and let $V_{L}$ be the graded vertex algebra over $\mathbb{Z}$ attached to $L$. Then, the functor that takes an affine scheme $\operatorname{Spec}R$ to the group of homogeneous $R$-vertex algebra automorphisms of $V_{L,R}$ is represented by a finite type smooth affine group scheme $\operatorname{Aut}V_{L}$ over $\mathbb{Z}$. This group is given as a product $G_{L}O(\tilde{L})$, where the normal subgroup $G_{L}$ is the split reductive group given in Definition 2.1.9, and $O(\tilde{L})$ is the finite flat group scheme defined in Proposition 3.2.2. The intersection of $G_{L}$ and $O(\tilde{L})$ is the Tits subgroup of the normalizer of the torus $T_{L}$ in $G_{L}$, i.e., the canonical elementary 2-torsion extension of the Weyl group. The quotient $\operatorname{Aut}V_{L}/G_{L}$ is isomorphic to the outer automorphism group $\operatorname{Aut}L/W_{L}$ of $L$. Furthermore, the functor that takes an affine scheme $\operatorname{Spec}R$, where $\det L$ is invertible in $R$, to the group of $R$-vertex operator algebra automorphisms of $V_{L,R}$ is represented by the base change $(\operatorname{Aut}V_{L})_{\mathbb{Z}[1/\det L]}$.
@@
The product property follows from the following sequence of reductions: By Corollary 4.2.3, $\operatorname{Aut}V_{L}=G_{L}N_{\operatorname{Aut}V_{L}}T_{L}$, but by combining Propositions 4.3.5 and 4.3.6, we have $N_{\operatorname{Aut}V_{L}}T_{L}=T_{L}O(\tilde{L})$, and the $T_{L}$ can be absorbed into the $G_{L}$.
We now show that the intersection of $G_{L}$ and $O(\tilde{L})$ is the Tits subgroup. By Lemma 4.3.10, $\tilde{W}_{L}$ is contained in the intersection, so it suffices to show the reverse inclusion. Since $O(\tilde{L})$ normalizes $T_{L}$, the intersection is contained in $N_{G_{L}}(T_{L})$. Taking the quotient by $T_{L}$, we find that the subgroup $O(\tilde{L})$ is mapped to $O(L)$, and the intersection is taken to $W_{L}$. Thus, the intersection is contained in the pullback of $W_{L}$ in $O(\tilde{L})$, which is precisely the Tits group by Lemma 4.3.10.
From the product property and the “second isomorphism theorem” for groups, the quotient $\operatorname{Aut}V_{L}/G_{L}=O(\tilde{L})G_{L}/G_{L}$ is isomorphic to the quotient of $O(\tilde{L})$ by the intersection $\tilde{W}_{L}$, and by Lemma 4.3.10, this is the outer automorphism group of $L$.

### v2

@@ 3.3. Action of orthogonal transformations on the lattice vertex algebra
The fact that this is a faithful group action follows from 3.3.2. The fact that the action is by vertex operator algebra automorphisms follows from the multiplication formulas for elements of $V_{L,R}$ - in the case $L$ is the Leech lattice and $R$ is a $\mathbb{Q}$-algebra, this is [Frenkel-Lepowsky-Meurman 1988] Theorem 10.4.6, but the proof does not depend on the particular lattice. When $R=\mathbb{Z}$, the formulas follow from base change to $\mathbb{Q}$, and the proof for general $R$ follows by base change from $\mathbb{Z}$. ∎
@@
The action of O(L~)O($\tilde${L}) on $V_{L}$ normalizes the $T_{L}$-action.
@@ 4.2. Local Conjugation
We wish to show that $\operatorname{Aut}V_{L}$ is a product of $G_{L}$ and O(L~)O($\tilde${L}). The following lemma lets us reduce this problem to more local questions.
@@
To show that $\operatorname{Aut}V_{L}$ is a product of $G_{L}$ and O(L~)O($\tilde${L}), it suffices to show that, given a commutative ring $S$ and an $S$-point $g$ of $\operatorname{Aut}V_{L}$, there is a faithfully flat cover $\{U_{i}\to\operatorname{Spec}S\}$ such that each $g|_{U_{i}}$ lies in the image of the product maps on $U_{i}$.
@@
By Proposition 4.1.2, $G_{L}$ is normal in $\operatorname{Aut}V_{L}$, so conjugation by $g$ takes $T_{L,S}$ to some torus $T^{\prime}$ in $G_{L,S}$. By [SGA3] Exp. XXII Corollary 5.3.11, any two maximal tori (of an fpqc locally split reductive group) are conjugate étale locally on $S$. Thus, there exists a choice of étale cover $\{S\to S_{i}\}$ and $\{g_{i}\in G_{L}(S_{i})\}$ so that conjugation by each $g_{i}$ takes $T^{\prime}_{S_{i}}$ to $T_{L,S_{i}}$. ∎
@@ 4.3. Normalizer of the distinguished torus
Consider the homomorphism $N_{\operatorname{Aut}V_{L}}T_{L}\to\operatorname{Aut}T_{L}$ induced by conjugation. We let $c:N_{\operatorname{Aut}V_{L}}T_{L}\to GL(L)$ be the composite of this homomorphism with the isomorphism $\operatorname{Aut}T_{L}\overset{\sim}{\to}GL(L)$.
@@
The orbit of $T_{L}$ acting on any nonzero vector in $\pi^{L}_{\lambda,R}$ is one dimensional, and contained in a line. Let $u\in\pi^{L}_{\lambda,R}$, and write $\sigma u=\sum_{i\in I}u_{i}$ for $u_{i}\in\pi^{L}_{\mu_{i},R}$, where the map $i\mapsto\mu_{i}$ is injective. If the indexing set $I$ has more than one element, then the orbit of $T_{L}$ on $\sigma u$ is not contained in a line. This contradicts the fact that $\sigma$ acts linearly and bijectively. We conclude that $I$ is a singleton. ∎
@@
The normalizer of $T_{L}$ in $\operatorname{Aut}V_{L}$ is the product $O(\tilde{L})C_{\operatorname{Aut}V_{L}}T_{L}$. In particular, $c$ induces an isomorphism $(N_{\operatorname{Aut}V_{L}}T_{L})/C_{\operatorname{Aut}V_{L}}T_{L}\overset{\sim}{\to}O(L)$ taking a normalizing automorphism to the induced orthogonal transformation on the lattice parametrizing Heisenberg modules in $V_{L}$.
@@
By Proposition 3.3.3, we know that $N_{\operatorname{Aut}V_{L}}T_{L}$ contains O(L~)O($\tilde${L}). Thus, it suffices to show $N_{\operatorname{Aut}V_{L}}T_{L}\subseteq O(\tilde{L})C_{\operatorname{Aut}V_{L}}T_{L}$.
The conjugation homomorphism $c:N_{\operatorname{Aut}V_{L}}T_{L}\to GL(L)$ has kernel $C_{\operatorname{Aut}V_{L}}T_{L}$ and by Lemma 4.3.4, its image lies in $O(L)$. The centralizer of $T_{L}$ in O(L~)O($\tilde${L}) is the kernel of the homomorphism to $O(L)$, and is therefore identified with $\operatorname{Hom}(L,\mu_{2})$. Thus, the image of O(L~)O($\tilde${L}) under $c$ is precisely $O(L)$, and we have an isomorphism. ∎
We can now show that $T_{L}$ is its own centralizer in the automorphism group scheme of $V_{L}$.
@@
It suffices to show that any homogeneous $R$-vertex algebra automorphism $\sigma$ of $V_{L,R}$ fixes $\omega$. By Corollary 4.2.3 and Proposition 4.3.5, we may write $\sigma$ as a product $gh$, where $g\in G_{L}(R^{\prime})$ and $h\in O(\tilde{L})(R^{\prime})$ for $R\hookrightarrow R^{\prime}$ étale. The formula for $\omega$ can be written as a sum over dual basis vectors, so it is invariant under orthogonal transformations, and in particular, the action of O(L~)O($\tilde${L}). It therefore suffices to show that $\omega$ is $G_{L}$-invariant, but for this it suffices to show that the action of $T_{L}$ and the infinitesimal actions of root vectors $h_{0}$ are trivial. Because $\omega$ lies in lattice-degree zero, $T_{L}$ acts trivially. For any root vector $h$, the well-known commutator formula $h_{m},L(n)]=mh_{m+n}$ ([Frenkel-Lepowsky-Meurman 1988] eq. 8.7.20) implies $h_{0}\omega=h_{0}L(-2)\mathbf{1}=L(-2)h_{0}\mathbf{1}=0$. ∎
We now consider the intersection of O(L~)O($\tilde${L}) and $G_{L}$ in $\operatorname{Aut}V_{L}$.
@@
Viewed as a subgroup of $N_{\operatorname{Aut}V_{L}}T_{L}$, $\tilde{W}_{L}$ is a normal subgroup of O(L~)O($\tilde${L}), given by the preimage of $W_{L}$ in $O(L)$ under the $\operatorname{Hom}(L,\mu_{2})$-cover $\pi:O(\tilde{L})\to O(L)$. In particular, the quotient $O(\tilde{L})/\tilde{W}_{L}$ is isomorphic to the outer automorphism group $O(L)/W_{L}$ of $L$.
@@
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(L,\mu_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tilde{W}_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${W_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${T_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{G_{L}}T_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${W_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${T_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\operatorname{Aut}V_{L}}T_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${O(L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
By Lemma 3.3.4, O(L~)O($\tilde${L}) normalizes $T_{L}$, i.e., we have an embedding $O(\tilde{L})\hookrightarrow N_{\operatorname{Aut}V_{L}}T_{L}$. By Proposition 3.2.2, restriction of $N_{\operatorname{Aut}V_{L}}T_{L}\to N_{\operatorname{Aut}V_{L}}T_{L}/T_{L}=O(L)$ to O(L~)O($\tilde${L}) is a surjective homomorphism to $O(L)$, with kernel $\operatorname{Hom}(L,\mu_{2})$. We therefore have a commutative diagram of extensions, where the vertical arrows are inclusions of normal subgroups, and the left square is a pullback:
${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Hom}(L,\mu_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${O(\tilde{L})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${O(L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1}$${1\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${T_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${N_{\operatorname{Aut}V_{L}}T_{L}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${O(L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${1.}$
By the universal property of pullback, this diagram fits into the pullback diagram along the embeddings $N_{G_{L}}T_{L}\to N_{\operatorname{Aut}V_{L}}T_{L}$ and $O(\tilde{L})\to N_{\operatorname{Aut}V_{L}}T_{L}$ as follows:
@@
where the vertical and diagonal arrows are inclusions of normal subgroups, and the quadrilaterals without equal signs are pullbacks. From the upper-right pullback rectangle, we obtain an isomorphism $O(\tilde{L})/\tilde{W}_{L}\simeq O(L)/W_{L}$. ∎
@@ 4.4. Main Theorem
[…]
@@
[…]
We now show that the intersection of $G_{L}$ and O(L~)O($\tilde${L}) is the Tits subgroup $\tilde{W}_{L}$. By Lemma 4.3.10, $\tilde{W}_{L}$ is contained in the intersection, so it suffices to show the reverse inclusion. Since O(L~)O($\tilde${L}) normalizes $T_{L}$, the intersection is contained in $N_{G_{L}}(T_{L})$. Taking the quotient by $T_{L}$, we find that the subgroup O(L~)O($\tilde${L}) is mapped to $O(L)$, and the intersection is taken to $W_{L}$. Thus, the intersection is contained in the pullback of $W_{L}$ in O(L~)O($\tilde${L}), which is precisely the Tits group by Lemma 4.3.10.
From the product property and the “second isomorphism theorem” for groups, the quotient $\operatorname{Aut}V_{L}/G_{L}=O(\tilde{L})G_{L}/G_{L}$ is isomorphic to the quotient of O(L~)O($\tilde${L}) by the intersection $\tilde{W}_{L}$, and by Lemma 4.3.10, this is the outer automorphism group of $L$.